OpenStudy (anonymous):

waht is the solution to the system ? -3x-y=-3 9x+3y=12

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

do i do it like inequalities?

jimthompson5910 (jim_thompson5910):

multiply everything in equation 1 by 3 to get -3x-y=-3 3*(-3x)-3*(y)=3*(-3) -9x - 3y = -9

jimthompson5910 (jim_thompson5910):

we now have -9x - 3y = -9 9x+3y=12

jimthompson5910 (jim_thompson5910):

add the two equations to get -9x-3y = -9 9x+3y = 12 ------------ 0x + 0y = 3 so we end up with 0 = 3

OpenStudy (anonymous):

3,3 ?

jimthompson5910 (jim_thompson5910):

is 0 = 3 true?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

0 is equal to 3? yes or no?

jimthompson5910 (jim_thompson5910):

0 is the same as 3?

OpenStudy (anonymous):

oh no

jimthompson5910 (jim_thompson5910):

0 = 3 is always false

jimthompson5910 (jim_thompson5910):

so there are no solutions because of that

OpenStudy (anonymous):

Awesome thanks

jimthompson5910 (jim_thompson5910):

yw